# The Uniform Functional Interpretation with Informative Types

### v1

@@ 1 Introduction
Functional interpretations of arithmetic were introduced by Kurt Gödel in his seminal dialectica paper [24]. Since then, the original interpretation has been refined, modified and studied in different aspects and for various purposes. It was also extended to analysis [44], admissible set theory [2, 13, 14] and to logic [18, 22, 23, 37]. A good reference, up to the end of the millennium, is the paper [1]. In the last three decades, Gödel’s interpretation in combination with majorizability has been successfully used as an applied tool for extracting computational information from ordinary proof of mathematics (specially in the domain of iterative arguments in nonlinear functional analysis). We are referring here to Ulrich Kohlenbach’s program of proof mining, conveniently explained in the book [31]. More recently, the book [9] appeared, authored by one of the early students of Gödel’s interpretation and covering material from a traditional perspective. It is not the aim of these brief lines to review the literature on functional interpretations, far too long and intricate, but rather to convey to the reader the idea that functional interpretations remain, to this day, an active area of study, with both theoretical and applied components.
The present paper rests on two main ideas. The first is that we regard functional interpretations as based on an ur interpretation, dubbed the uniform interpretation, whose main feature is that quantifications are not computationally relevant. The second idea is that computational relevance is acquired via so-called type-informative predicates. Being an element of a type is contingent upon falling under a given type-informative predicate, whose interpretation is computationally relevant. Let us briefly discuss these two aspects in turn.
Uniform quantifications already appear in the works of Jean-Louis Krivine [35] and Ulrich Berger [4], but within a realizability framework. If we write $\langle A\rangle^{\bm{a}}$ for saying that witness $\bm{a}$ realizes $A$, then the quantifier clauses for uniform quantification are:
@@
The second idea of the paper is, roughly, saying that being an element of a certain type is like falling under a given primitive unary relation symbol, and that the interpretations of these (so-called, type-informative) predicate symbols may be computationally relevant. For the purposes of this introduction, we denote by $x:\sigma$ the unary predicate saying that $x$ is of type $\sigma$ (we use a different notation in the body of the text). For instance, we can have $\langle x:\mathbb{N}\rangle^{n}$ as being $x\leq n$. This is the so-called bounding interpretation at the ground type $\mathbb{N}$ (Section 8). Alternatively, we could have $\langle x:\mathbb{N}\rangle^{n}$ as being $x=n$, the precise interpretation at the ground type $\mathbb{N}$ (Section 7).
@@ 2.2 The uniform interpretation and its soundness
Let $\bm{r}^{\prime},\bm{s}$ and $\bm{p}$ be terms so that we have the equalities $\bm{r}^{\prime}\bm{a}\bm{c}\bm{v}=\bm{r}\bm{a}\bm{c}\bm{c}\bm{v}$, $\bm{p}\bm{a}\bm{c}=\bm{q}\bm{a}\bm{c}\bm{c}$ and $\bm{s}\bm{a}\bm{c}\bm{v}=\bm{t}\bm{a}\bm{c}\bm{c}\bm{v}\cup\bm{t}^{\prime}\bm{a}\bm{c}\bm{c}\bm{v}$. This choice clearly works. It is, of course, the possibility of accumulating counter-witnesses (by taking a union) in the definition of the term $\bm{s}$ that necessitates the Diller-Nahm interpretation (vis à vis the dialectica interpretation).
@@ 3 Relativization and type-informative interpretations
We will consider different base interpretations of ${\mathcal{L}}_{{\mathbf{S}}}^{\mathcal{I}}$. The following three interpretations of ${\mathcal{I}}_{\sigma}(x^{\sigma})$ will be considered in this paper.
@@
It is clear that the sentences of the above two conditions imply ${\mathcal{I}}_{\sigma}(t)$, for every closed term $t$ of type $\sigma$. Since our framework assumes that, for each base type $X$, there is a constant $e_{X}$ of type $X$ then, from these constants and lambda-abstraction (combinators), it is well-known that we can recursively define a closed term for each type of the source theory. As a consequence, the type-informative predicates are non-empty. This observation is important because, in the next subsection, we define a relativization of formulas to the type-informative predicates and, for the relativization to operate properly, these predicates must be non-empty.
@@ 3.1 The main soundness theorem
The $\mathcal{I}$}-interpretation of $\forall\bm{x}^{\bm{\sigma}}A(\bm{x}\bm{)}$ asks for closed terms $\bm{t}$ such that, the target theory $\mathbf{T}$} proves
@@
where the proof takes place in $\mathbf{S}$$_{{\mathcal{I}}}$. The axioms $(i)$, $(ii)$ and $(iii)$ of $\mathbf{S}$$_{{\mathcal{I}}}$ are $\mathcal{U}$}-interpretable by our assumption that we are given a $\mathbf{T}$}-sound type-informative base interpretation of $\mathbf{S}$}. Therefore, by the Uniform Soundness Theorem (see Theorem 2.9), there are tuples of closed witnessing terms $\bm{q},\bm{s}$ and $\bm{t}$ such that such that $\mathbf{T}$} proves the following: for all $\bm{a}$, $\bm{c}$ and $\bm{d}$ and for all $\bm{x}^{\bm{\sigma}}$,
@@ 3.2 Some refinements
This is the interpretation of the conclusion of the rule. The quantifier-free rule of extensionality due Gödel [25] and Spector [44] is an example of an admissible rule as described above.
@@ 5 The canonical interpretation of the function types
The following lemma is fundamental for dealing with canonical interpretations that interpret disjunction uniformly:
@@ 6 Interlude on arithmetic
We should also say a few words regarding the target theory $\mathbf{T}$} where the interpretations of the theory $\mathbf{HA}^{\omega}$. We focus on describing its terms (amongst which are the witnessing terms, with an associated term reduction calculus). For the concrete examples of the next sections, the language of the target theory is the language of $\mathbf{HA}^{\omega}$, endowed with the star types and the vocabulary of the star combinatory calculus (including recursors pertaining to the star types) together with a witnessing constant $\mathrm{m}$ of type $\mathbb{N}^{*}\to\mathbb{N}$ whose intended meaning is that $\mathrm{m}F$, or $\mathrm{m}(F)$, is the maximum of the non-empty finite set $F$, where $F\subseteq\mathbb{N}$. Regarding the target theories proper (i.e., their axioms), as commented in Section 2, we do not need to be too precise about them.
@@
This amounts to taking the target theory $\mathbf{T}$} as the set of truths in $\mathcal{S}^{\omega}_{*}$ (formulated the language of the target theory). The restriction of $\mathcal{S}^{\omega}_{*}$ to the types built without the star-construct is denoted by $\mathcal{S}^{\omega}$. As it is well-known, the arithmetical part of $\mathcal{S}^{\omega}$ is a model of Gödel’s theory $\mathsf{T}$.
@@ 7.1 Canonical upon the precise interpretation ${\mathcal{I}}^{\mathrm{cp}}$
The ${\mathcal{I}}^{\mathrm{cp}}$-interpretation is the $\mathcal{I}$}-interpretation obtained from the base interpretation which is the precise interpretation at the ground types $\mathbb{N}$ and $\mathbb{B}$, and the canonical interpretation at function types.
@@ 8.1 Majorizability upon the bounding interpretation ${\mathcal{I}}^{\mathrm{mb}}$
The ${\mathcal{I}}^{\mathrm{mb}}$-interpretation is the $\mathcal{I}$}-interpretation obtained from the base interpretation of arithmetic that defines the type-informative predicates as follows:
@@ 8.3 Canonical upon the bounding interpretation ${\mathcal{I}}^{\mathrm{cb}}$
The ${\mathcal{I}}^{\mathrm{cb}}$-interpretation is the base interpretation of arithmetic that defines the type-informative predicates as follows: it is the bounding interpretation at the ground type $\mathbb{N}$, the uniform interpretation at the ground type $\mathbb{B}$, and it is extended canonically at the function types.
@@
It is clear that one has $\langle{\mathcal{I}}_{\mathbb{N}}({\mathsf{0}})\rangle^{0}$ and $\forall k\langle{\mathcal{I}}_{\mathbb{N}\to\mathbb{N}}({\mathsf{S}})\rangle^{S}_{k}$, where $S$ is the successor function. With the following result, we show that the ${\mathcal{I}}^{\mathrm{cb}}$-interpretation is a type-informative base interpretation of $\mathbf{HA}^{\omega}$ (the proposition below is similar to Proposition 7.4, but the proof is not quite the same):
@@
The proof is similar to that of Proposition 7.4. Our aim is to show that
@@ 9 On the uniformity of restricted quantifiers
For a formula $B(x^{\sigma},\bm{y})$, with a distinguished free-variable $x^{\sigma}$, we can introduce in the source theory $\mathbf{S}$} primitive restricted (or bounded) quantifiers $\forall x\unlhd_{\sigma}B(\bm{y})\,A(x)$ and $\exists x\unlhd_{\sigma}B(\bm{y})\,A(x)$, with axioms
$\forall x\unlhd_{\sigma}B(\bm{y})\,A(x)$ $\;\leftrightarrow\;\forall x^{\sigma}(B(x,\bm{y})\to A(x))$ (14)
$\exists x\unlhd_{\sigma}B(\bm{y})\,A(x)$ $\;\leftrightarrow\;\exists x^{\sigma}(B(x,\bm{y})\wedge A(x)).$ (15)
@@
The above example works because we are using the canonical upon the bounding interpretation ${\mathcal{I}}^{\mathrm{cb}}$. Were we using instead the majorizability upon the bounding interpretation ${\mathcal{I}}^{\mathrm{mb}}$, then the $\mathcal{U}$}-interpretation of (16) would ask for closed terms $t^{\mathbb{N}\to\mathbb{N}}$ and $q^{\mathbb{N}^{*}}$, such that
@@
Suppose that the source theory $\mathbf{S}$} is extended to include primitive $B$-restricted quantifiers with axioms (15) and (14). Let $\mathbf{S}$$_{{\mathcal{I}}}$ be the auxiliary theory of the proof of Theorem 3.20 together with (16) as a new axiom. Then, we can extend the relativization procedure of Definition 3.11 as follows
  • $\begin{array}[]{lcl}(\forall x\unlhd_{\sigma}B(\bm{y})\,A(x))^{\mathcal{I}}&:\equiv&\forall x^{\sigma}(B(x^{\sigma},\bm{y})^{\mathcal{I}}\to A(x)^{\mathcal{I}})\\[5.69054pt] (\exists x\unlhd_{\sigma}B(\bm{y})\,A(x))^{\mathcal{I}}&:\equiv&\exists x^{\sigma}(B(x^{\sigma},\bm{y})^{\mathcal{I}}\wedge A(x)^{\mathcal{I}}).\end{array}$
@@
It is enough to see that, in the presence of (16), the relativizations of (15) and (14) are logically trivial. Consider the case of $\forall x\unlhd_{\sigma}B(\bm{y})\,A(x)$. By definition, its relativization is
@@
In other words, whenever we can $\mathcal{U}$}-interpret that $B(x^{\sigma},\bm{y})$ is “stronger” than ${\mathcal{I}}_{\sigma}(x)$, we can simplify the $\mathcal{I}$}-relativization of the $B$-restricted quantifiers. Hence, simpler $\mathcal{I}$}-interpretations of $\exists x\unlhd_{\sigma}B(\bm{y})\,A(x)$ and $\forall x\unlhd_{\sigma}B(\bm{y})\,A(x)$ are obtained. Let us formalize this by extending Proposition 3.14.
@@
$\begin{array}[]{lcl}|\forall x\unlhd_{\tau}B(\bm{y})\,A(x)|^{\bm{f},\bm{g}}_{\bm{u},\bm{b}}&\mathrm{\,is\,}&\forall x^{\sigma}(\forall\bm{v}\in\bm{g}\bm{u}\bm{b}\,|B(x,\bm{y})|^{\bm{u}}_{\bm{v}}\,\to|A(x)|^{\bm{f}\bm{u}}_{\bm{b}})\\[5.69054pt] |\exists x\unlhd_{\tau}B(\bm{y})A(x)|^{\bm{u},\bm{a}}_{\bm{V},\bm{U}}&\mathrm{\,is\,}&\exists x^{\sigma}(\forall\bm{v}\in\bm{V}\,|B(x,\bm{y})|^{\bm{u}}_{\bm{v}}\wedge\forall\bm{b}\in\bm{U}\,|A(x)|^{\bm{a}}_{\bm{b}}).\end{array}$

### v2

@@ 1 Introduction
Functional interpretations of arithmetic were introduced by Kurt Gödel in his seminal dialectica paper [24]. Since then, the original interpretation has been refined, modified and studied in different aspects and for various purposes. It was also extended to analysis [44], admissible set theory [2, 13, 14] and to logic [18, 22, 23, 37]. A good reference, up to the end of the millennium, is the paper [1]. In the last three decades, Gödel’s interpretation in combination with majorizability has been successfully used as an applied tool for extracting computational information from ordinary mathematical proofs (specially in the domain of iterative arguments in nonlinear functional analysis). We are referring here to Ulrich Kohlenbach’s program of proof mining, conveniently explained in the book [32]. More recently, the book [9] appeared, authored by one of the early students of Gödel’s interpretation and covering material from a traditional perspective. It is not the aim of these brief lines to review the literature on functional interpretations, far too long and intricate, but rather to convey to the reader the idea that functional interpretations remain, to this day, an active area of study, with both theoretical and applied components.
The present paper rests on two main ideas. The first is that we regard functional interpretations as based on a background interpretation, dubbed the uniform interpretation, whose main feature is that quantifications are not computationally relevant. The second idea is that computational relevance is acquired via so-called type-informative predicates. Being an element of a type is contingent upon falling under a given type-informative predicate, whose interpretation is computationally relevant. Let us briefly discuss these two aspects in turn.
Uniform quantifications already appear in the works of Jean-Louis Krivine [35] and Ulrich Berger [4], but within a realizability framework. Let us write $\langle A\rangle^{\bm{a}}$ for saying that witness $\bm{a}$ realizes $A$. The quantifier clauses for uniform quantification are:
@@
The second idea of the paper is, roughly, saying that being an element of a certain type is like falling under a given primitive unary relation symbol, and that the interpretations of these (so-called, type-informative) predicate symbols may be computationally relevant. For the purposes of this introduction, we denote by $x:\sigma$ the unary predicate saying that $x$ is of type $\sigma$ (in the body of the text, we use a different notation ${\mathcal{I}}_{\sigma}(x)$ for the type-informative predicates). For instance, we can have $\langle x:\mathbb{N}\rangle^{n}$ as being $x\leq n$. This is the so-called bounding interpretation at the ground type $\mathbb{N}$ (Section 8). Alternatively, we could have $\langle x:\mathbb{N}\rangle^{n}$ as being $x=n$, the precise interpretation at the ground type $\mathbb{N}$ (Section 7).
@@ 2.2 The uniform interpretation and its soundness
Let $\bm{r}^{\prime},\bm{s}$ and $\bm{p}$ be terms so that we have the equalities $\bm{r}^{\prime}\bm{a}\bm{c}\bm{v}=\bm{r}\bm{a}\bm{c}\bm{c}\bm{v}$, $\bm{p}\bm{a}\bm{c}=\bm{q}\bm{a}\bm{c}\bm{c}$ and $\bm{s}\bm{a}\bm{c}\bm{v}=\bm{t}\bm{a}\bm{c}\bm{c}\bm{v}\cup\bm{t}^{\prime}\bm{a}\bm{c}\bm{c}\bm{v}$. This choice clearly works. It is, of course, the possibility of accumulating counter-witnesses (by taking a union) in the definition of the term $\bm{s}$ that necessitates the Diller-Nahm interpretation (vis-à-vis the dialectica interpretation).
@@ 3 Relativization and type-informative interpretations
We will consider different base interpretations of ${\mathcal{L}}_{{\mathbf{S}}}^{\mathcal{I}}$.
@@
It is clear that the sentences of the above two conditions imply ${\mathcal{I}}_{\sigma}(t)$, for every closed term $t$ of type $\sigma$. Since our framework assumes that, for each base type $X$, there is a constant $e_{X}$ of type $X$ then, from these constants and lambda-abstraction (combinators), it is well known that we can recursively define a closed term for each type of the source theory. As a consequence, the type-informative predicates are non-empty. This observation is important because, in the next subsection, we define a relativization of formulas to the type-informative predicates and, for the relativization to operate properly, these predicates must be non-empty.
@@ 3.1 The main soundness theorem
The $\mathcal{I}$}-interpretation of $\forall\bm{x}^{\bm{\sigma}}A(\bm{x})$ asks for closed terms $\bm{t}$ such that, the target theory $\mathbf{T}$} proves
@@
where the proof takes place in $\mathbf{S}$$_{{\mathcal{I}}}$. The axioms $(i)$, $(ii)$ and $(iii)$ of $\mathbf{S}$$_{{\mathcal{I}}}$ are $\mathcal{U}$}-interpretable by our assumption that we are given a $\mathbf{T}$}-sound type-informative base interpretation of $\mathbf{S}$}. Therefore, by the Uniform Soundness Theorem (see Theorem 2.9), there are tuples of closed witnessing terms $\bm{q},\bm{s}$ and $\bm{t}$ such that $\mathbf{T}$} proves the following: for all $\bm{a}$, $\bm{c}$ and $\bm{d}$ and for all $\bm{x}^{\bm{\sigma}}$,
@@ 3.2 Some refinements
This is the interpretation of the conclusion of the rule. The quantifier-free rule of extensionality due to Gödel [25] and Spector [44] is an example of an admissible rule as described above.
@@ 5 The canonical interpretation of the function types
The following proposition is fundamental for dealing with canonical interpretations that interpret disjunction uniformly:
@@ 6 Interlude on arithmetic
We should also say a few words regarding the target theory $\mathbf{T}$} where the interpretations of the theory $\mathbf{HA}^{\omega}$ takes place. We focus on describing its terms (among which are the witnessing terms, with an associated term reduction calculus). For the concrete examples of the next sections, the language of the target theory is the language of $\mathbf{HA}^{\omega}$, endowed with the star types and the vocabulary of the star combinatory calculus (including recursors pertaining to the star types) together with a witnessing constant $\mathrm{m}$ of type $\mathbb{N}^{*}\to\mathbb{N}$ whose intended meaning is that $\mathrm{m}F$, or $\mathrm{m}(F)$, is the maximum of the non-empty finite set $F$, where $F\subseteq\mathbb{N}$. Regarding the target theories proper (i.e., their axioms), as commented in Section 2, we do not need to be too precise about them.
@@
This amounts to taking the target theory $\mathbf{T}$} as the set of truths in $\mathcal{S}^{\omega}_{*}$ (formulated the language of the target theory). The restriction of $\mathcal{S}^{\omega}_{*}$ to the types built without the star-construct is denoted by $\mathcal{S}^{\omega}$. As it is well known, the arithmetical part of $\mathcal{S}^{\omega}$ is a model of Gödel’s theory $\mathsf{T}$.
@@ 7.1 Canonical upon the precise interpretation ${\mathcal{I}}^{\mathrm{cp}}$
The ${\mathcal{I}}^{\mathrm{cp}}$-interpretation is the base interpretation of arithmetic that defines the type-informative predicates as follows: it is precise at the ground types $\mathbb{N}$ and $\mathbb{B}$ and it is extended canonically at the function types.
@@ 8.1 Majorizability upon the bounding interpretation ${\mathcal{I}}^{\mathrm{mb}}$
The ${\mathcal{I}}^{\mathrm{mb}}$-interpretation is the base interpretation of arithmetic that defines the type-informative predicates as follows:
@@ 8.3 Canonical upon the bounding interpretation ${\mathcal{I}}^{\mathrm{cb}}$
The ${\mathcal{I}}^{\mathrm{cp}}$-interpretation is the base interpretation of arithmetic that defines the type-informative predicates as follows: it is the bounding interpretation at the ground type $\mathbb{N}$, the uniform interpretation at the ground type $\mathbb{B}$, and it is extended canonically at the function types.
@@
It is clear that one has $\langle{\mathcal{I}}_{\mathbb{N}}({\mathsf{0}})\rangle^{0}$ and $\forall k\langle{\mathcal{I}}_{\mathbb{N}\to\mathbb{N}}({\mathsf{S}})\rangle^{S}_{k}$, where $S$ is the successor function. With the following result, we show that the ${\mathcal{I}}^{\mathrm{cb}}$-interpretation is a type-informative base interpretation of $\mathbf{HA}^{\omega}$:
@@
The proof is similar to that of Proposition 7.4, with some important differences. Our aim is to show that
@@ 9 On the uniformity of restricted quantifiers
For a formula $B(x^{\sigma},\bm{y})$, with a distinguished free-variable $x^{\sigma}$, we can introduce in the source theory $\mathbf{S}$} primitive restricted (or bounded) quantifiers $\forall x\triangleleft_{\sigma}B(\bm{y})\,A(x)$ and $\exists x\triangleleft_{\sigma}B(\bm{y})\,A(x)$, with axioms
$\forall x\triangleleft_{\sigma}B(\bm{y})\,A(x)$ $\;\leftrightarrow\;\forall x^{\sigma}(B(x,\bm{y})\to A(x))$ (14)
$\exists x\triangleleft_{\sigma}B(\bm{y})\,A(x)$ $\;\leftrightarrow\;\exists x^{\sigma}(B(x,\bm{y})\wedge A(x)),$ (15)
where $A(x)$ can have parameters, including from $\bm{y}$.
@@
The above example works because we are using the canonical upon the bounding interpretation ${\mathcal{I}}^{\mathrm{cb}}$. If we were using instead the majorizability upon the bounding interpretation ${\mathcal{I}}^{\mathrm{mb}}$, then the $\mathcal{U}$}-interpretation of (16) would ask for closed terms $t^{\mathbb{N}\to\mathbb{N}}$ and $q^{\mathbb{N}^{*}}$, such that
@@
Suppose that the source theory $\mathbf{S}$} is extended to include primitive $B$-restricted quantifiers with axioms (15) and (14). Let $\mathbf{S}$$_{{\mathcal{I}}}$ be the auxiliary theory of the proof of Theorem 3.20 together with (16) as a new axiom. Then, we can extend the relativization procedure of Definition 3.11 as follows:
  • $\begin{array}[]{lcl}(\forall x\triangleleft_{\sigma}B(\bm{y})\,A(x))^{\mathcal{I}}&:\equiv&\forall x^{\sigma}(B(x^{\sigma},\bm{y})^{\mathcal{I}}\to A(x)^{\mathcal{I}})\\[5.69054pt] (\exists x\triangleleft_{\sigma}B(\bm{y})\,A(x))^{\mathcal{I}}&:\equiv&\exists x^{\sigma}(B(x^{\sigma},\bm{y})^{\mathcal{I}}\wedge A(x)^{\mathcal{I}}).\end{array}$
@@
It is enough to see that, in the presence of (16), the relativizations of (15) and (14) are logically trivial. Consider the case of $\forall x\triangleleft_{\sigma}B(\bm{y})\,A(x)$. By definition, its relativization is
@@
As we will see, whenever we can $\mathcal{U}$}-interpret that $B(x^{\sigma},\bm{y})$ is “stronger” than ${\mathcal{I}}_{\sigma}(x)$, we can simplify the $\mathcal{I}$}-relativization of the $B$-restricted quantifiers. Hence, simpler $\mathcal{I}$}-interpretations of $\exists x\triangleleft_{\sigma}B(\bm{y})\,A(x)$ and $\forall x\triangleleft_{\sigma}B(\bm{y})\,A(x)$ are obtained. Let us formalize this by extending Proposition 3.14.
@@
$\begin{array}[]{lcl}|\forall x\triangleleft_{\sigma}B(\bm{y})\,A(x)|^{\bm{f},\bm{g}}_{\bm{u},\bm{b}}&\mathrm{\,is\,}&\forall x^{\sigma}(\forall\bm{v}\in\bm{g}\bm{u}\bm{b}\,|B(x,\bm{y})|^{\bm{u}}_{\bm{v}}\,\to|A(x)|^{\bm{f}\bm{u}}_{\bm{b}})\\[5.69054pt] |\exists x\triangleleft_{\sigma}B(\bm{y})A(x)|^{\bm{u},\bm{a}}_{\bm{V},\bm{U}}&\mathrm{\,is\,}&\exists x^{\sigma}(\forall\bm{v}\in\bm{V}\,|B(x,\bm{y})|^{\bm{u}}_{\bm{v}}\wedge\forall\bm{b}\in\bm{U}\,|A(x)|^{\bm{a}}_{\bm{b}}).\end{array}$